OpenStudy (anonymous):

@ the beginning of a walk roberto and juanita are 5.5 miles apart. they leave at same time and walk in the same direction, roberto overtakes jaunita in 11 hours. if they walk towards each other they meet in 1 hour. what are their speeds?

OpenStudy (anonymous):

1(x+y)=5.5 11(x-y)=5.5 ----> x-y=0.5 (x+y)+(x-y)=5.5+0.5 2x=6 x=3 (roberto's speed) 5.5-3 = 2.5(Juanita's speed)